OpenStudy (anonymous):

Two cars leave town going in opposite directions. One car is traveling 60 miles per hour (mph), and the other car is traveling 50 mph. How long will it take before the cars are 220 miles apart?

OpenStudy (stormfire1):

Well, without doing any algebra you can conclude that after 2 hours, car 1 has gone 120 miles and car 2 has gone 100 miles...which totals 220 miles.

OpenStudy (stormfire1):

Just in case, here's how to do it algebraically:\[(x * 50)+(x * 60) = 220\] and just solve for x.

OpenStudy (anonymous):

2 hours 3 hours 4 hours 5 hours

OpenStudy (stormfire1):

Read my first post, the answer is there.

OpenStudy (anonymous):

Since the cars are moving in opposite direction then their resultant speed = (60+50) mph = 110 mph Time = Distance/ speed Distance = 220 miles speed = 110 mph so time=distance/speed = 220/110 = 2 hours.